OpenStudy (anonymous):

Solve: 3x+2y=15 2x-10=y

OpenStudy (anonymous):

\[\Large \begin{array}{l} 3x + 2y = 15\\ 2x - 10 = y\\ 3x + 2(2x - 10) = 15\\ 3x + 4x - 20 = 15\\ 7x = 35\\ x = \frac{{35}}{7} = 5\\ y = 2 \cdot 5 - 10 = 0 \end{array}\]

OpenStudy (anonymous):

fewscrewsmissing thats the biggest font ive ever seen in open study

OpenStudy (anonymous):

Big is best.

OpenStudy (anonymous):

(5, 0)